OpenStudy (dtan5457):

Isaac has written down three integers. The sum of all 3 when cubed is 7624. The sum of the range of the numbers (smallest and largest) is 27. The middle integer is 13. What are the other integers?

OpenStudy (welshfella):

(x + y + z)^3 = 7624 y = 13 x + z = 27

OpenStudy (dtan5457):

i have that im trying to solve x^3+z^3=7455 x+y=27

OpenStudy (dtan5457):

*x+z=27

OpenStudy (welshfella):

u sure you have the question right? 7624 is not a perfect cube

OpenStudy (dtan5457):

i think they mean the sum of each cube

OpenStudy (mathstudent55):

@welshfella I think you need to do x^3 + y^3 + z^3 = 7624

OpenStudy (welshfella):

x^3 + y^3 + z^3 = 7624

OpenStudy (welshfella):

ok

OpenStudy (dtan5457):

yea

OpenStudy (anonymous):

Isaac has written down three integers. The sum of all 3 when cubed is 7624. The sum of the range of the numbers (smallest and largest) is 27. The middle integer is 13. What are the other integers? WLOG assume \(a\le b <c\) $$a^3+b^3+c^3=7624\\a+c=27\\b=13\\\implies a+b+c=40$$ now observe $$(a+b+c)^3=a^3+b^3+c^3+3(a^2b+ab^2+a^2c+ac^2+b^2c+bc^2)+6abc\\40^3=7624+3(13a^2+169a+a^2c+ac^2+169c+13c^2)+78ac$$

OpenStudy (welshfella):

as a first step I would try x = 12 and z = 15 and x = 11 and z = 16

OpenStudy (welshfella):

trial and error

OpenStudy (dtan5457):

is it possible to solve algebraically?

OpenStudy (mathstudent55):

x^3 + 13^3 + z^3 = 7624 x^3 + 2197 + z^3 = 7624 x^3 + z^3 = 5427 x + z = 27

OpenStudy (welshfella):

weel you can generate an equation in x by plugging in x = 27 - z

OpenStudy (anonymous):

now since \(a+c=27\) we get \((a+c)^2=a^2+c^2+2ac\implies a^2+c^2=27^2-2ac\) so consider now $$13a^2+169a+169c+13c^2+a^2c+ac^2+78ac\\\quad =169(a+c)+13(a^2+c^2)+ac(a+c+78)$$

OpenStudy (welshfella):

sorry z = 27 - x

OpenStudy (welshfella):

x^3 + (27-x)^3 = 7455

OpenStudy (anonymous):

oops, that should be $$3(13a^2+169a+169c+13c^2+a^2c+ac^2)+78ac$$ so $$3\cdot (169(a+c)+13(a^2+c^2)+ac(a+c+26))$$

OpenStudy (anonymous):

we have that \(a+c=27\implies a+c+26=53\) and further that \(a^2+c^2=729-2ac\) so $$64000=7624+3\cdot(169\cdot27+13\cdot(729-2ac)+ac\cdot53)$$

OpenStudy (dtan5457):

Never mind guys, i solved this using some quadratic formula. thanks for your help though

OpenStudy (anonymous):

so $$\frac{64000-7624}3-169\cdot27-13\cdot729=27ac\\\implies ac=176$$

OpenStudy (dtan5457):

I just did x^3+13^3+z^3=7624 x^3+z^3=5427 x+z=27 x=(-z+27) (-z+27)^3+z^3=5427 -z^3+81z^2-2187z+27^3+z^3=5427 81z^2-2187z+14256=0 ^^quadratic formula and you get the other 2 integers

OpenStudy (mathstudent55):

x = 27 - z x^3 = (27 - z)^3 x^3 = -z^3 + 81z^2 - 2187z + 19683 -z^3 + 81z^2 - 2187z + 19683 + z^3 = 5427 81z^2 - 2187z + 19683 = 5427 81z^2 - 2187z + 14256 = 0 z^2 - 27z + 176 = 0 (z - 11)(z - 16) = 0 z = 11 or z = 16

OpenStudy (welshfella):

i8 tried that and got -27x^2 as first term - must have made a mistake somewhere

OpenStudy (welshfella):

yea i see where i went wrong

OpenStudy (anonymous):

now we know $$ac=176\\a+c=27$$ so$$a(27-a)=176\\27a-a^2=176\\a^2-27a+176=0\\a=\frac{27\pm\sqrt{729-2\cdot176}}2\\\implies a=11\\\implies c=16$$

OpenStudy (anonymous):

but yeah, you can plug in for \(b\) at the start and get $$a^3+c^3=\dots\\a^3+c^3=(a+c)^3-3ac(a+c)$$ and get \(ac\) much faster

OpenStudy (anonymous):

thing is, once you have \(a+c,ac\), writing out the quadratic and factoring is unnecessary because you ask the same question -- what two numbers add up to \(27\) but multiply to give \(176\)?

OpenStudy (dtan5457):

you are right but i guess its just a little easier to go with things we learned before sometimes

OpenStudy (dtan5457):

even if it takes longer